# PATCH! Psychometrics-AssisTed benCHmarking of Large Language Models: A Case Study of Proficiency in 8th Grade Mathematics

### v2

@@ 6 Conclusions
In this paper, we propose PATCH, a psychometrics-inspired framework to address current limitations of LLM benchmarks, especially for the purpose of model and human comparison. We demonstrate PATCH with an 8th grade mathematics proficiency test, where PATCH yields evaluation outcomes that diverge from those based on existing benchmarking practices. This underscores the potential of PATCH to reshape the LLM benchmarking landscape. Nevertheless, our paper has several limitations. First, PATCH requires validated tests, which can be resource-intensive if tests need to be developed from scratch. However, this also opens up opportunities for collaboration between LLM researchers, psychometricians and test institutes. Second, the validity, reliability, and fairness of using tests validated solely on humans for LLM benchmarking are debatable due to possibly differing notions of proficiency and cognitive processes between LLMs and humans. Nonetheless, such tests are still better than non-validated benchmarks, particularly for comparison of model and human performance. Advancing LLM benchmarking further requires tests validated on LLMs (and humans for model-human comparisons), necessitating theoretical work on LLM-specific constructs and the development of LLM-specific IRT models and testing procedures. Third, our experiment only includes two proprietary LLMs and one proficiency test. We consider this sufficient for demonstrating PATCH, but not enough if the goal is to benchmark all relevant LLMs across different tests.
@@ Acknowledgements
We thank Anna Wegmann, Yupei Du, Melody Sepahpour-Fard, Elise Herrewijnen, Gianluca Sperduti for their helpful suggestions and comments. This work was supported by the Dutch Research Council (NWO) (grant number VI.Vidi.195.152 to D. L. Oberski; grant number VI.Veni.192.130 to D. Nguyen).
@@ Appendix B Example Responses
Note that the example responses below correspond to the example test items in Appendix A.1.
@@
Example 3 (prompt variant 1, GPT4-V, correct)
Type of Question: [Free Text Response] This is a free-text response question.
To find the value of $x$, we will use the formula for the volume of a rectangular box, which is:
@@
$200=x\times 5\times 4$
To solve for $x$, we divide both sides of the equation by $5\times 4$:
$x=\frac{200}{5\times 4}$
$x=\frac{200}{20}$
@@
Answer: $x=10\,\text{cm}$
Appendix C Detailed Result Figures
Figure 3: Distribution of proficiency estimates for GPT-4V, Gemini-Pro-Vision and all participants of TIMSS 2011 8th grade mathematics test. Left figure (A) shows the proficiency estimates based on the percentages of correct responses. Right figure (B) shows the IRT-based proficiency estimates. The middle vertical line in each box plot represents the weighted mean proficiency score, with the error bars indicating its 95% confidence interval. The borders of each box indicate the range of the middle 50% of all values, with the two whiskers indicating the 5th and 95th percentiles.

### v3

@@ 6 Conclusion
Our paper has the following limitations, among others. First, PATCH requires validated tests, which can be resource-intensive if tests need to be developed from scratch. However, this also opens up opportunities for collaboration between LLM researchers, psychometricians and test institutes. Second, the validity, reliability, and fairness of using tests validated solely on humans for LLM benchmarking are debatable due to possibly differing notions of proficiency and cognitive processes between LLMs and humans. Nonetheless, such tests are still better than non-validated benchmarks, particularly for comparison of model and human performance. Advancing LLM benchmarking further requires tests validated on LLMs (and humans for model-human comparisons), necessitating theoretical work on LLM-specific constructs and the development of LLM-specific IRT models and testing procedures. Third, our experiment only includes four LLMs and one proficiency test. We consider this sufficient for demonstrating PATCH, but not enough if the goal is to benchmark as many LLMs as possible across different tests.
@@ Acknowledgements
We thank our reviewers, Anna Wegmann, Yupei Du, Melody Sepahpour-Fard, Elise Herrewijnen, Gianluca Sperduti for their helpful suggestions and comments. This work was supported by the Dutch Research Council (NWO) (grant number VI.Vidi.195.152 to D. L. Oberski; grant number VI.Veni.192.130 to D. Nguyen).
@@ Appendix B Example Responses
Note that the example responses below correspond to (the order of) the example test items in Appendix A.1.
@@
Example 3 (prompt variant 1, Qwen-VL-Max, correct)
Is it drawing, free text response, or multiple choice? It’s a multiple-choice question.
To find the volume of a rectangular prism (box), we use the formula:
@@
200 $200\text{ }\text{cm}^{3}=x\times 5\text{ cm}\times 4\text{ cm}$
Now, let’s solve for $x$:
$x=\frac{200\text{ }\text{cm}^{3}}{5\text{ cm}\times 4\text{ cm}}$
$x=\frac{200\text{ }\text{cm}^{3}}{20\text{ cm}^{2}}$
@@
[…]
Appendix C TIMSS Datasets
TIMSS 2011 Mathematics Eighth Grade
This dataset was used in this study to demonstrate the PATCH framework. See Section 5.1 for details.
Source: User Guide, Items and International Database for TIMSS 2011: Science – Eighth Grade. Copyright ©2013 International Association for the Evaluation of Educational Achievement (IEA). Publisher: TIMSS & PIRLS International Study Center, Lynch School of Education, Boston College.
Our study contributes three additional datasets. Similar to the dataset above, they are also based on officially released items by TIMSS but differ in the test subject, school grade level and/or test year. We constructed each dataset by using a mix of manual labour and OCR tools to extract item details from the official PDFs of the released items. The resulting dataset consists of a LaTeX file ("main.tex") and a folder of item-related images. The test items are formatted in LLM-friendly format. With these three additional datasets, we hope to facilitate interested researchers to benchmark LLMs using these datasets with our PATCH framework. See below for more detail.
TIMSS 2011 Mathematics Fourth Grade
This dataset is similar to the one we used to demonstrate PATCH but focuses on a different fourth grade mathematics with 73 items covering three domains: number, geometric shape and measures, and data display. It can be used to benchmark LLMs against representative samples of fourth-grade students from 57 regions.
Source: User Guide, Items and International Database for TIMSS 2011: Mathematics – Fourth Grade. Copyright ©2013 International Association for the Evaluation of Educational Achievement (IEA). Publisher: TIMSS & PIRLS International Study Center, Lynch School of Education, Boston College.
TIMSS 2008 Advanced Mathematics
This dataset focuses on assessing proficiency in advanced mathematics at the end of secondary high school. It can be used to benchmark LLMs against representative samples of final-year students in secondary school from 10 countries who have taken an advanced mathematics course. There are 40 items in total, covering algebra, calculus and geometry.
Source: TIMSS Advanced 2008 User Guide and Items for the International Database: Advanced Mathematics. Copyright ©2009 International Association for the Evaluation of Educational Achievement (IEA). Publisher: TIMSS & PIRLS International Study Center, Lynch School of Education, Boston College.
TIMSS 2008 Advanced Physics
This dataset focuses on assessing proficiency in advanced physics at the end of secondary high school. It can be used to benchmark LLMs against representative samples of final-year students in secondary school from 10 countries who have taken an advanced physics course. There are 39 items in total, covering mechanics, atomic and nuclear physics, electricity and magnetism, as well as heat and temperature.
Source: TIMSS Advanced 2008 User Guide and Items for the International Database: Advanced Physics. Copyright ©2009 International Association for the Evaluation of Educational Achievement (IEA). Publisher: TIMSS & PIRLS International Study Center, Lynch School of Education, Boston College.
Licences
According to the website of TIMSS 2011131313https://timssandpirls.bc.edu/timss2011/international-database.html and 2008141414https://timssandpirls.bc.edu/timss_advanced/idb.html:
TIMSS and PIRLS are registered trademarks of IEA. Use of these trademarks without permission of IEA by others may constitute trademark infringement. Furthermore, the website and its contents, together with all online and/or printed publications and released items by TIMSS, PIRLS, and IEA are and will remain the copyright of IEA.
All publications and released items by TIMSS, PIRLS, and IEA, as well as translations thereof, are for non-commercial, educational, and research purposes only. Prior notice is required when using IEA data sources or datasets for assessments or learning materials. IEA reserves the right to refuse copy deemed inappropriate or not properly sourced.
Therefore, our use of TIMSS data in this research is in accordance with the intended use.
Appendix D Use of AI Assistants
We used ChatGPT to improve the writing of limited parts of the paper. We also used Mathpix to perform OCR on the PDFs containing the TIMSS released items before further processing into appropriate format. No AI was used for coding or analyses.
Appendix E Detailed Result Figure
See next page.
Figure 3: Distribution of proficiency estimates for GPT-4V, Gemini-Vision-Pro, Qwen-VL-Plus, Qwen-VL-Max and all participating regions of TIMSS 2011 8th grade mathematics test. Left figure (A) shows the proficiency estimates based on the percentages of correct responses. Right figure (B) shows the IRT-based proficiency estimates. The middle vertical line in each box plot represents the weighted mean proficiency score, with the error bars indicating its 95% confidence interval. The borders of each box indicate the range of the middle 50% of all values, with the two whiskers indicating the 5th and 95th percentiles. Note that we adhere to the official naming conventions of TIMSS 2011 when reporting the names of participating regions, with no intent to offend anyone.